OpenStudy (anonymous):

e^x>1+x+(x^2)/2 (x>0) prove the inequalities.

OpenStudy (zzr0ck3r):

are you looking at power series in class?

OpenStudy (anonymous):

do you know the taylor series expansion of e^x?

mathslover (mathslover):

(First prove that : \(e^x=1+x+(x^2)/2!+(x^3)/3!+... \) by using power series) i.e. \(e^x = 1 + x +\cfrac{x^2}{2} + (\text{ something positive } ) \) Thus \(e^x > 1 + x + \cfrac{x^2 }{ 2} \)

mathslover (mathslover):

Try to prove the given equation : \(e^x = 1 + x + \cfrac{x^2 }{ 2! } + \cfrac{x^3 }{ 3!} + ... \)

OpenStudy (anonymous):

@ni21 have you already learned about Taylor series?

OpenStudy (anonymous):

dear all, actually i have the answer but i don't understand.. in the solution , it shows using f' and f''... does this related to taylor series?

OpenStudy (anonymous):

related, yes, but it would be hard to explain that way... how about post the answer and then I can explain it for you

OpenStudy (anonymous):

f' and f'' are related to taylor series

OpenStudy (anonymous):

try looking at 4.2 and 4.3 here http://www.math.smith.edu/~rhaas/m114-00/chp4taylor.pdf

OpenStudy (anonymous):

and just plug in 0 for a

OpenStudy (anonymous):

\[f(x)=e^x-[1+x+(x^2)/2]\]\[f'(x)=e^x-1-x\]\[f''(x)=e^x-1>e^0-1=0\]\[(x>0)\] from f''(x)>0, f'(x)>f'(0) since \[f'(0)=e^0-1-0=0, f(x)>0\] from f'(x)>0, f(x)>f(0) since f(0)=e^0-(1+0+0)=0, f(x)>0 therefore., \[e^x>1+x+(x^2)/2 when x>0\]

OpenStudy (anonymous):

oh. It is defining a function as the difference between e^x and 1+x+(x^2)/2, then finding the behavior of that function by finding the first and second derivatives.

OpenStudy (anonymous):

so if f(x) is positive, the statement is true.

OpenStudy (anonymous):

@Peter14 ... does this mean the solution given is using mathematical induction? i mean how can i know that i need to find until f'' or f'?

OpenStudy (anonymous):

there is another question that actually find up to f' only.. here is the question.. \[\ln (1+x)>\frac{ x }{ 1+x} (x>0)\] prove the inequalities

OpenStudy (anonymous):

the solution doesn't use mathematical induction, or at least not that I can see.

OpenStudy (anonymous):

if that so, then how can i know the solution needs f' or f''? million thanks in advance, @Peter14:) other members, i need your opinions, too:)

OpenStudy (anonymous):

here. http://en.wikipedia.org/wiki/Racetrack_principle do you see how this applies?

OpenStudy (anonymous):

i think i can see how it links... but if let say i stuck in the middle, can i ask you again?:)

OpenStudy (anonymous):

sure, i'll be here

OpenStudy (anonymous):

so for this question, ln(1+x)>x 1+x (x>0) prove the inequalities.. the method is same right?

OpenStudy (anonymous):

yes, the method is the same.

OpenStudy (anonymous):

thanks,@Peter14

OpenStudy (anonymous):

hi @Peter14 , i try to solve for ln(1+x)>x 1+x (x>0), i mean try to prove this inequality but when i check the solution, it shows that i dont have to find f''. so now, how do i know whether i need to find up to f'' or f'? is there any condition or am i missed out anything that is related to?

OpenStudy (anonymous):

let's do it together...

OpenStudy (anonymous):

what's f'(x)?

OpenStudy (anonymous):

(sorry, my latex isn't working)

OpenStudy (anonymous):

\[\frac{ x }{ (1+x)^2 } \]

OpenStudy (anonymous):

i simplified from this \[\frac{ 1 }{ 1+x }-\frac{ 1 }{ (1+x)^2 }\] to \frac{ x }{ (1+x)^2 }

OpenStudy (anonymous):

so when x is positive, is x/((1+x)^2) always positive?

OpenStudy (anonymous):

yes. f'(x) always positive

OpenStudy (anonymous):

so then you don't need to find f''(x) because you can see from this that the derivative of ln (1+x) > the derivative of the other thing

OpenStudy (anonymous):

if we were to look back at the first question that i've posted, when \[f'(x)=e^x-1-x\] means it is still not enough to prove that it is always positive? is it?

OpenStudy (anonymous):

it's because e^x is sorta hard to evaluate without a calculator, so it's hard to prove just from that that e^x>1+x, when you go to f''(x) it's easy because you know e^x is always greater than 1 for x>0

OpenStudy (anonymous):

@ni21 you see?

OpenStudy (anonymous):

i'm still trying to digest :) so to put in other example, if i manage to f'(x)=sinx-xcosx, i still cannot prove that it is always positive and need to derive 1 more time, right?

OpenStudy (anonymous):

yes, I think so

OpenStudy (anonymous):

what was f(x) for that?

OpenStudy (anonymous):

-2 cos x - x sin x ?

OpenStudy (anonymous):

\[2(1-cosx)-xsinx\]

OpenStudy (anonymous):

more or less the same.. without constant 2 yet the basic is still there:) that's impressive!

OpenStudy (anonymous):

that's indefinite integrals

OpenStudy (anonymous):

except I left out the C as is my common mistake

OpenStudy (anonymous):

but please note that the condition for x i given as such (0<x<=pi/2)

OpenStudy (anonymous):

ok, that's nice to know

OpenStudy (anonymous):

yes, take the derivative of sin x - x cos x and all will be clear

OpenStudy (anonymous):

glad to know that i'm in the right path. just to double confirm with regarding the reason of f'(x) and f"(x); we only need to derive up to f''(x) because to assure that when x is positive then the derivative is always positive, am i right? or is there any concrete reason?

OpenStudy (anonymous):

yes, you're right, though the english there is a bit awkward.

OpenStudy (anonymous):

haha:) i know my english really poor.

OpenStudy (anonymous):

it works, it's enough to get the message across, and I hadn't noticed before now.

OpenStudy (anonymous):

you mean my english?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

anything else?

OpenStudy (anonymous):

that should be enough for this time. arigatou gozaimasu! nice to learn from you.

OpenStudy (anonymous):

nice to teach you